find the exact value of cos 2 theta given cosine of theta = negative 12 over 13 and theta is between 180 degrees and 270 degrees.
Answer by stanbon(75887)

find the exact value of cos 2 theta given cosine of theta = negative 12 over 13 and theta is between 180 degrees and 270 degrees.
------
cos(2t) = cos^(t)-sin^2(t) = 2cos^2(t)-1
---------------------------------
cos(t) = -12/13
-----
Therefore: cos(2t) = 2(-12/13)^2-1 = 0.7041
-------------------
Cheers,
Stan H.
-----------------